OpenStudy (anonymous):

Joselyn is a manager at a sign painting company. She has three painters, Allen, Brianne, and Charles. Allen can complete a large project in 16 hours. Brianne can complete the same sized project in 18 hours. Charles is new, so no one knows how long it will take him. Joselyn assigns them all a large project to complete together. Explain to Joselyn how this project can tell her how long it would take Charles if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

allens rate is \(\frac{1}{16}\) and brianne's is \(\frac{1}{18}\) if charles can do the work in \(t\) hours his rate is \(\frac{1}{t}\)